# VSearcher: Long-Horizon Multimodal Search Agent via Reinforcement Learning

### 03 Mar 2026 v1

@@ Abstract
Large models are increasingly becoming autonomous agents that interact with real-world environments and use external tools to augment their static capabilities. However, most recent progress has focused on text-only large language models, which are limited to a single modality and therefore have narrower application scenarios. On the other hand, multimodal large models, while offering stronger perceptual capabilities, remain limited to static knowledge and lack the ability to access and leverage up-to-date web information. In this paper, we propose VSearcher, turning static multimodal model into multimodal search agent capable of long-horizon, multi-turn tool use in real-world web environments, including text search, image search, and web browsing, via reinforcement learning. Specifically, we introduce Iterative Injection Data Synthesis pipeline to generate large-scale, complex multimodal QA questions, which are further filtered with comprehensive metrics to ensure high quality and sufficient difficulty. We then adopt an SFT-then-RL training pipeline to turn base multimodal models to agent capable of multi-turn tool calling in real-world web environments. Besides, we propose a multimodal search benchmark $M$-SearchExam dedicated to evaluating search capabilities of multimodal search agents, which proves highly challenging for recent proprietary models. Extensive evaluations across multiple multimodal search benchmarks reveal effectiveness of our method. VSearcher achieves superior performance compared to recent multimodal search agents and even surpasses several proprietary models on multimodal web search tasks.
@@ 1 Introduction
  • Challenging benchmark for evaluation of multimodal search agent. We curate a challenging benchmark $M$-SearchExam for evaluating multimodal browsing ability, on which strong proprietary models achieve low performance.
@@ 2 Method
In this section, we provide an overview of our post-training framework for VSearcher, including Iterative Injection-based Data Synthesis, Rejection Sampling Fine-tuning, and Reinforcement Learning. We also describe the curation details of our proposed multimodal search benchmark, $M$-SearchExam.
@@ 2.2 Rejection Sampling Finetuning
where NN is the number of training trajectories, $o_{i,t}$ denotes the $t$-th token of the output sequence for the $i$-th trajectory, $T_{i}$ is the total length of the output sequence of the $i$-th training trajectory, $o_{i,<t}$ represents the tokens preceding the $t$-th token of the $i$-th training trajectory, $q_{i}$ is the input query for the $i$-th training trajectory, and $\pi_{\theta}$ is the model policy parameterized by $\theta$.
@@ 2.3 Reinforcement Learning
2.4 Curation of $M$-SearchExam
To enable more thorough evaluation of multimodal search agents, we curate a challenging multimodal search benchmark, $M$-SearchExam. $M$-SearchExam is constructed with our proposed Iterative Injection-based Data Synthesis pipeline. Specifically, we perform 10 rounds of iterative extension to make the browsing tasks highly challenging. We also apply data filtering procedure to ensure high quality and difficulty. Additionally, we explicitly remove seeds that appear in our training data to avoid data leakage. In total, $M$-SearchExam contains 283 synthesized browsing tasks. Quantitative experiments show that several proprietary models achieve low performance on $M$-SearchExam, highlighting difficulty of the benchmark and the effectiveness of our Iterative Injection-based Data Synthesis pipeline in generating highly challenging browsing tasks.
@@ 6 Conclusion
In this paper, we propose VSearcher, a long-horizon, multi-turn, tool-using multimodal search agent trained with reinforcement learning in real-world web environment. VSearcher can adaptively leverage text search, image search, and visit tools to retrieve web information and solve challenging tasks. We propose a systematic post-training framework to convert static multimodal model into multimodal search agent. In particular, we propose Iterative Injection-based Data Synthesis pipeline, together with rigorous data filtering strategies, to automatically synthesize highly challenging and high-quality multimodal browsing tasks at scale. We further adopt rejection-sampling fine-tuning to transfer strong multi-turn browsing ability from proprietary teacher model to base multimodal models, and apply reinforcement learning to further generalize the tool-use capabilities. For more thorough evaluation of multimodal search agent, we curate a highly challenging benchmark, $M$-SearchExam, on which several recent proprietary models yield low performance. Across extensive evaluations on multiple multimodal search benchmarks, VSearcher consistently outperforms recent multimodal search agents and even surpasses several proprietary models.
@@ A.5 Benchmark
$M$-BrowseComp. $M$-BrowseComp contains 224 challenging, hand-crafted questions designed to evaluate multimodal retrieval ability. The questions are constructed in order to make sure text alone is insufficient to answer correctly. We filter out samples without image input, resulting in 122 samples for evaluation.

### 06 Mar 2026 v2

@@ Abstract
Large models are increasingly becoming autonomous agents that interact with real-world environments and use external tools to augment their static capabilities. However, most recent progress has focused on text-only large language models, which are limited to a single modality and therefore have narrower application scenarios. On the other hand, multimodal large models, while offering stronger perceptual capabilities, remain limited to static knowledge and lack the ability to access and leverage up-to-date web information. In this paper, we propose VSearcher, turning static multimodal model into multimodal search agent capable of long-horizon, multi-turn tool use in real-world web environments, including text search, image search, and web browsing, via reinforcement learning. Specifically, we introduce Iterative Injection Data Synthesis pipeline to generate large-scale, complex multimodal QA questions, which are further filtered with comprehensive metrics to ensure high quality and sufficient difficulty. We then adopt an SFT-then-RL training pipeline to turn base multimodal models to agent capable of multi-turn tool calling in real-world web environments. Besides, we propose a multimodal search benchmark MM-SearchExam dedicated to evaluating search capabilities of multimodal search agents, which proves highly challenging for recent proprietary models. Extensive evaluations across multiple multimodal search benchmarks reveal effectiveness of our method. VSearcher achieves superior performance compared to recent multimodal search agents and even surpasses several proprietary models on multimodal web search tasks.
@@ 1 Introduction
  • Challenging benchmark for evaluation of multimodal search agent. We curate a challenging benchmark MM-SearchExam for evaluating multimodal browsing ability, on which strong proprietary models achieve low performance.
@@ 2 Method
In this section, we provide an overview of our post-training framework for VSearcher, including Iterative Injection-based Data Synthesis, Rejection Sampling Fine-tuning, and Reinforcement Learning. We also describe the curation details of our proposed multimodal search benchmark, MM-SearchExam.
@@ 2.2 Rejection Sampling Finetuning
where $N$ is the number of training trajectories, $o_{i,t}$ denotes the $t$-th token of the output sequence for the $i$-th trajectory, $T_{i}$ is the total length of the output sequence of the $i$-th training trajectory, $o_{i,<t}$ represents the tokens preceding the $t$-th token of the $i$-th training trajectory, $q_{i}$ is the input query for the $i$-th training trajectory, and $\pi_{\theta}$ is the model policy parameterized by $\theta$.
@@ 2.3 Reinforcement Learning
2.4 Curation of MM-SearchExam
To enable more thorough evaluation of multimodal search agents, we curate a challenging multimodal search benchmark, MM-SearchExam. MM-SearchExam is constructed with our proposed Iterative Injection-based Data Synthesis pipeline. Specifically, we perform 10 rounds of iterative extension to make the browsing tasks highly challenging. We also apply data filtering procedure to ensure high quality and difficulty. Additionally, we explicitly remove seeds that appear in our training data to avoid data leakage. In total, MM-SearchExam contains 283 synthesized browsing tasks. Quantitative experiments show that several proprietary models achieve low performance on MM-SearchExam, highlighting difficulty of the benchmark and the effectiveness of our Iterative Injection-based Data Synthesis pipeline in generating highly challenging browsing tasks.
@@ 6 Conclusion
In this paper, we propose VSearcher, a long-horizon, multi-turn, tool-using multimodal search agent trained with reinforcement learning in real-world web environment. VSearcher can adaptively leverage text search, image search, and visit tools to retrieve web information and solve challenging tasks. We propose a systematic post-training framework to convert static multimodal model into multimodal search agent. In particular, we propose Iterative Injection-based Data Synthesis pipeline, together with rigorous data filtering strategies, to automatically synthesize highly challenging and high-quality multimodal browsing tasks at scale. We further adopt rejection-sampling fine-tuning to transfer strong multi-turn browsing ability from proprietary teacher model to base multimodal models, and apply reinforcement learning to further generalize the tool-use capabilities. For more thorough evaluation of multimodal search agent, we curate a highly challenging benchmark, MM-SearchExam, on which several recent proprietary models yield low performance. Across extensive evaluations on multiple multimodal search benchmarks, VSearcher consistently outperforms recent multimodal search agents and even surpasses several proprietary models.
@@ A.5 Benchmark
MM-BrowseComp. MM-BrowseComp contains 224 challenging, hand-crafted questions designed to evaluate multimodal retrieval ability. The questions are constructed in order to make sure text alone is insufficient to answer correctly. We filter out samples without image input, resulting in 122 samples for evaluation.